OpenStudy (dls):

Which of the following functions of time represent A)Periodic motion B)Non periodic motion Give the time period in each case (w is any +ve constant...)

OpenStudy (dls):

\[\LARGE \sin wt+\cos wt\]

OpenStudy (dls):

\[\LARGE \sin wt+\cos2wt+\sin4wt\]

OpenStudy (dls):

\[\LARGE e^{-wt} \]

OpenStudy (dls):

and log(wt)

OpenStudy (yrelhan4):

*

OpenStudy (dls):

Did the * mean you don't know or you tagged yourself ?O_o

OpenStudy (dls):

@Mashy @shubhamsrg

OpenStudy (shubhamsrg):

sinwt + sin(pi/2 -wt) 2 sin(pi/4) cos(pi/2 -2wt) 2sin(pi.4) sin(2w)t T = 2pi/2w

OpenStudy (shubhamsrg):

I applied the formula wrong hmm

OpenStudy (shubhamsrg):

it'll be 2sin(pi/4) cos(pi/4 - wt)

OpenStudy (shubhamsrg):

T = 2pi/w

OpenStudy (anonymous):

1 ,2 periodic.. 3 ,4 non periodic :p